Shiftsum for OWL 9/13/99

Alexander Ledovskoy (sasha@uvahea.phys.virginia.edu)
Mon, 13 Sep 1999 08:01:49 -0400 (EDT)

Shift Summary for OWL Shift, 09/13/99

On Shift: Elizabeth, Sasha L.

Hours of beam available at all intensities: 4.17h
Hours of beam available at nominal intensity: 4.17h
Hours of E832 data (special runs not included): 4.17h
Hours of special runs (laser scans, muon runs...): ~1.0h Laser scan

Accounting of lost time: none

========================================================================

Runs ended this shift:

14509 BEAM_832 23.4 M events
14511 Laser_scan

Tapes brought to FCC this shift:


========================================================================

Accesses during shift:

========================================================================

Overview of shift: 832 data taking at 4.0 E12
2 h 50 min lost due to TeV quench (we used this time
for laser scan and tape change)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

0:24 TeV quench
Tapes 93% full we decided to change tapes

Beam estimate 2 hours
14511 0:46 Laser scan
1:33 Laser scan is finished, still no beam

14512 1:47 BEAM_832 run started, still no beam

68 3:15 Beam returns

375 DA_DEAD=3.75 error

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

If humidity alarms go off (humidity above 10% in the blockhouse)
call Leonard Nelson. His home number is 630-231-1416; pager
888-750-0702.